OpenStudy (anonymous):

Need help with integral: I know I am supposed how I am supposed to use partial fractions, but the proffesor did it, I think she got it wrong, I want to confirm, but maybe I am wrong (x-8) / (x^3 -4x^2 + 4x) dx

OpenStudy (amistre64):

may have to use partial fraction decomp

OpenStudy (amistre64):

the denom factors to x(x-2)^2

OpenStudy (anonymous):

then you put them A/ X + B/(X-2) + C/(X-2)^2 RIGHT?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

x-8 = A(x-2)(x-2) + Bx(x-2) + Cx when x = 0 -8 = A(4); A = -2 ................................. when x = 2 -6 = C(2); C = -3 and so on

OpenStudy (anonymous):

Ooh ok I think I got it from here. Than you :)

OpenStudy (amistre64):

youre welcome :)